OpenStudy (anonymous):

How to evaluate (3^-3+3^-4)/3^-5

OpenStudy (anonymous):

\[\frac{3^{-3}+3^{-4}}{3^{-5}}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

probably simplest to break it into two fractions \[\frac{3^{-3}}{3^{-5}}+\frac{3^{-4}}{x^{-5}}\]

OpenStudy (anonymous):

Where x is 3.

OpenStudy (anonymous):

oh, okay let me try breaking it up

OpenStudy (anonymous):

So it's 12. Thanks!

OpenStudy (anonymous):

when you divide if the bases are the same you subtract the exponents. be careful when you subtract, because you are subtracting a Negative number. \[\frac{3^{-3}}{3^{-5}}=3^{-3-(-5)}=3^{-3+5}=3^2\]

OpenStudy (anonymous):

12 is correct. Yes.

OpenStudy (anonymous):

yes it is 12. good work

OpenStudy (anonymous):

hello polpak!

OpenStudy (anonymous):

\[\color{green}{\text{hello polpak}}\]

OpenStudy (anonymous):

Hello there.